OpenStudy (anonymous):

which equation can help you determine whether a triangle with side lengths of 3, 4, and 5 is a right triangle? MEDAL GRANTED

OpenStudy (amoodarya):

 \[a^2=b^2+c^2\]

OpenStudy (anonymous):

the choices im given are: 3 to the second power + 5 to the second power = 4 to the second power or

OpenStudy (amoodarya):

no 5>4 5>3 so \[5^2=4^2+3^2=\\25=16 +9\]

OpenStudy (anonymous):

or (3+4) second power = 5 to second power

OpenStudy (amoodarya):

no 3^2=9 4^2=16 3 2nd power +4 2nd power =9+16=25 and 25 is 5 to the second power

OpenStudy (amoodarya):